# Integrating large language models and active inference to understand eye movements in reading and dyslexia

### v2

@@ Abstract
We present a novel computational model employing hierarchical active inference to simulate reading and eye movements. The model characterizes linguistic processing as inference over a hierarchical generative model, facilitating predictions and inferences at various levels of granularity, from syllables to sentences. Our approach combines the strengths of large language models for realistic textual predictions and active inference for guiding eye movements to informative textual information, enabling the testing of predictions. The model exhibits proficiency in reading both known and unknown words and sentences, adhering to the distinction between lexical and nonlexical routes in dual route theories of reading. Notably, our model permits the exploration of maladaptive inference effects on eye movements during reading, such as in dyslexia. To simulate this condition, we attenuate the contribution of priors during the reading process, leading to incorrect inferences and a more fragmented reading style, characterized by a greater number of shorter saccades. This alignment with empirical findings regarding eye movements in dyslexic individuals highlights the model’s potential to aid in understanding the cognitive processes underlying reading and eye movements, as well as how reading deficits associated with dyslexia may emerge from maladaptive predictive processing. In summary, our model represents a significant advancement in comprehending the intricate cognitive processes involved in reading and eye movements, with potential implications for understanding and addressing dyslexia through the simulation of maladaptive inference. It may offer valuable insights into this condition and contribute to the development of more effective interventions for treatment.
keywords:
@@
Hierarchical active inference, Predictive coding, Large Language Models, Reading, Dyslexia
[…]
[CNR]organization=Institute of Cognitive Sciences and Technologies - National Research Council
,addressline=Via San Martino della Battaglia 44, city=Rome, postcode=00185, state=Italy, country=Italy
Introduction
Processing natural language – encompassing understanding, reading, and producing linguistic content – represents a fundamental ability of our species. Extensive research in psychology, neuroscience, linguistics, and machine learning has explored the intricate ways we process natural language.
@@ Introduction
Neuroscientific studies have revealed that natural language processing is inherently hierarchical, involving multiple brain regions and the integration of various sensory inputs (Capek et al., 2004; Butterfuss and Kendeou, 2018; Price, 2012). This hierarchical processing spans from individual letters, phonemes, and words to complete sentence comprehension, with a hierarchy of brain areas actively maintaining these elements in working memory across multiple time scales (Hasson et al., 2015).
@@
Recent studies have increasingly supported the idea of hierarchical predictive coding as a formal theory describing perception as an inferential process, involving reciprocal exchanges between predictions and prediction errors across brain hierarchies (Rao and Ballard, 1999; Friston, 2005). In particular, various computational neuroimaging studies employing large language models (LLMs) during linguistic tasks have provided compelling evidence for both the predictive nature of language processing and the prediction hierarchies proposed by predictive coding (Goldstein et al., 2022; Schmitt et al., 2021; Baldassano et al., 2017; Zadbood et al., 2017; Chang et al., 2022; Caucheteux et al., 2023; Heilbron et al., 2022).
A recent computational neuroimaging study investigated human electrocorticographic (ECoG) responses to narratives using LLMs trained to predict the next word. Remarkably, the study revealed that like the LLMs, the brain engages in next-word prediction before word onset, computes prediction error signals, and utilizes latent representations of words (embeddings) contextualized based on the sequence of prior words (Goldstein et al., 2022).
@@
Converging evidence emerges from three recent fMRI studies utilizing deep learning models during linguistic tasks. These studies not only confirm the predictive nature of language processing but also lend support to the prediction hierarchies proposed by predictive coding. The first fMRI study (Schmitt et al., 2021) trareferined a LLM to predict the next word at multiple time scales, identifying event boundaries as high surprise (also explored in (Baldassano et al., 2017; Zadbood et al., 2017; Chang et al., 2022)). Analyzing human functional magnetic resonance data during story listening, the study revealed an event-based hierarchy of surprise signals evolving along temporoparietal regions, with surprise signals gating bottom-up and top-down connectivity across neighboring time scales.
Furthermore, another fMRI study provided compelling evidence of the brain’s ability to predict a hierarchy of representations spanning multiple timescales in the future (Caucheteux et al., 2023). Enhancing LLMs with the capability to predict beyond the next word increased their fit with human data. The study also highlighted the hierarchical organization of brain predictions, ranging from temporal cortices predicting shorter-range representations (e.g., the next word) to frontoparietal cortices predicting higher-level, longer-range, and more contextual representations.
@@
Notably, a separate fMRI study reported that evoked brain responses to words are influenced by linguistic predictions and a metric of unexpectedness, closely aligning with the hierarchical scheme of active inference, where lower-level predictions are informed by higher-level predictions (Heilbron et al., 2022). Additionally, the study demonstrated that these hierarchical predictions can be well-aligned with standard levels of analysis in psycholinguistics, including meaning, grammar, words, and speech sounds, reinforcing the validity of the standard decomposition.
Taken together, these studies, alongside others (Kutas and Hillyard, 1984; Weissbart et al., 2020; Koskinen et al., 2020; Willems et al., 2016; Frank et al., 2015), provide compelling support for the significance of prediction and hierarchical predictive coding in language processing. Moreover, they underscore the growing relevance of LLMs in comprehending linguistic processing (Goldstein et al., 2022).
@@
Despite these advancements, the above studies have primarily focused on LLMs that passively receive sensory information, rather than actively searching for it. However, linguistic tasks, such as reading written text and listening to speech, are inherently active processes (Ferro et al., 2010; Norris, 2006; Friston et al., 2021). For example, during reading, eye movements (saccades) actively guide attention to relevant parts of the text, rather than processing every piece of text linearly. This active reading process suggests that saccades play a crucial role in hypothesis testing, selecting informative parts of the text to test predictions (Ferro et al., 2010; Norris, 2006; Donnarumma et al., 2017; Friston et al., 2012).
Moreover, eye movement studies during reading have proven to be valuable in understanding reading difficulties, such as dyslexia. Dyslexia, a common reading disorder affecting 5-10% of the population, is associated not only with atypical brain activation patterns during language processing (Price, 2012) but also with atypical eye movement patterns, such as increased numbers of forward saccades and decreased saccade lengths compared to control groups (Zoccolotti et al., 2005; De Luca et al., 1999; Lyon et al., 2003; Franzen et al., 2021).
The connection between prediction-based language processing and the perceptual challenges observed in dyslexia has been explored in previous studies like (Jaffe-Dax et al., 2015). These investigations suggest that dyslexia could be characterized as a computational difficulty in effectively combining prior information, such as implicit memory of previous words, with noisy observations like the currently perceived word. Essentially, when prior information is given excessively low weighting relative to its internal noise, the time required to form a coherent understanding of text (or to perform auditory discrimination, as in (Jaffe-Dax et al., 2015)) significantly increases because novel observations consistently surprise dyslexics. Through computational simulations, the authors demonstrated that this pattern of results could account for dyslexics’ perceptual deficit. However, the proposed model lacks the capability to actively select information, such as generating eye movements.
@@
Several reading models incorporating eye movements, such as E-Z Reader (Reichle et al., 1998), SWIFT (Engbert et al., 2005; Seelig et al., 2020; Rabe et al., 2021), Über-Reader (Reichle, 2021; Veldre et al., 2020), Glenmore (Reilly and Radach, 2002, 2006), and OB1-Reader (Snell et al., 2018) significantly advanced our understanding of the role of reading dynamics. These models have successfully replicated numerous empirical findings, highlighting how word-level attributes like length, frequency, and predictability impact reading dynamics. However, these models focus largely on data regression and they do not provide a mechanistic explanation of the cognitive processes integral to reading, such as for example how saccades are generated.
@@
In this paper, we propose a novel computational model that unifies hierarchical predictive processing and hypothesis testing during reading by integrating the LLM BERT (Bidirectional Encoder Representations from Transformers) (Devlin et al., 2019) with active inference (Friston, 2010; Parr et al., 2022). Our model views reading as an active (Bayesian) inference problem, employing a hierarchical generative model to represent causal relationships between textual elements at different levels (letters, syllables, words, and sentences). By generating predictions at each level and testing them through saccades, our model actively simulates reading.
@@
The model incorporates three significant insights discussed earlier. Firstly, it conceptualizes linguistic processing as inference, employing a hierarchical generative model that allows for predicting and inferring at different timescales, such as syllables, words, and sentences. By integrating the LLM BERT (Devlin et al., 2019) at the highest hierarchical level, our model can handle realistic reading tasks effectively, processing sentences of arbitrary length. Additionally, thanks to its hierarchical structure, our model can read text both word-by-word, a common approach in LLMs (i.e., the lexical route), and syllable-by-syllable, akin to the nonlexical route in dual route theories of reading (Coltheart, 2005).
@@
Secondly, the model utilizes its generative capabilities not only for language recognition and prediction, similar to LLMs, but also for simulating eye movements and saccades. In our model, reading involves an active, hypothesis testing process: the model generates saccades to the most informative parts of the text to validate its predictions and disambiguate among competing hypotheses about the content being read (Ferro et al., 2010; Baldassano et al., 2017; Donnarumma et al., 2017; Friston et al., 2012; Norris, 2006).
Thirdly, by reducing the influence of prior information in the inference (Jaffe-Dax et al., 2015), our model produces simulated eye movements during reading that align with empirical findings in dyslexia (Zoccolotti et al., 2005; De Luca et al., 1999; Hutzler and Wimmer, 2004; Rayner, 1983). This aspect allows the model to generate eye movement patterns consistent with dyslexic individuals’ reading impairments.
This model allows us to gain deeper insights into the integration of information during reading, the role of saccades in testing predictions, and the potential impact of maladaptive predictive processing in reading deficits associated with dyslexia.
Furthermore, its hierarchical composition from primitive content unveils phenomena challenging to investigate with other models, such as reading unfamiliar or unknown words.
@@ 1 Methods
In this section, we provide a summary description of the model, a brief introduction to active inference (Parr et al., 2022), and a detailed discussion of the hierarchical active inference reading model used in this paper.
@@ 1.1 Brief explanation of the model
We present a novel hierarchical active inference model for reading and eye movements during reading (Friston, 2010). Consistent with evidence of hierarchical language processing in the brain (Heilbron et al., 2022), our model consists of three levels representing syllables, words, and sentences, as illustrated in Figure 1(a).
@@
At each level, the inference process involves integrating three types of messages: bottom-up messages from nodes at the level below (conveying observations), top-down messages from nodes at the level above (conveying predictions), and lateral messages from nodes representing the previous timestep at the same level (providing memory). For instance, the inference about the current syllable is informed by bottom-up observations about the currently observed letter, top-down predictions from the currently inferred word, and lateral information about the previous syllable. Similarly, the inference about the current word relies on bottom-up observations (i.e., the currently estimated word), top-down predictions from the level above (i.e., at the sentence level), and information about the previous word.
@@
Figure 1: The hierarchical active inference model of reading and eye movements. (a) The hierarchical model comprises three layers for syllables (Level 1), words (Level 2), and sentences (Level 3), represented using the formalism of Partially Observable Markov Decision Processes (POMDPs). Empty nodes indicate hidden variables, filled nodes represent observations, and edges depict probabilistic relations between the nodes, as detailed in B. (b) Model parameters. (c) The sequence of saccades generated by the model while reading an example sentence. Yellow dots indicate the positions of the saccades in the text, and colored lines trace the sequence of saccades from red to yellow. (d) Evolution of the probabilities of correctly recognized syllables (blue lines), words (orange lines), and sentences (yellow lines) over time. The blue circles at the bottom level indicate saccades. Once the first level confidently infers a syllable (or when the maximum level of iterations is reached), it sends a bottom-up signal to the second level (indicated by the orange vertical bar) to aid in word inference. Similarly, the second level sends a signal to the third level (indicated by the yellow vertical bar) upon confidently inferring a word. Please refer to the main text for further explanation.
To ensure that our model is capable of reading realistic text, we utilize the large language model BERT (Devlin et al., 2019) to generate a probability distribution over the 100 most likely sentences consisting of 4-8 words. Importantly, this process can be iterated to read sentences of varying lengths, providing the flexibility to handle texts of different complexities.
@@
Lastly, our model has the capability to determine where to make (forward or backward) saccades and select the next letter observation (Donnarumma et al., 2017). All three levels of the model are involved in the process of selecting where to direct the next saccade, by selecting the word (Level 3), syllable (Level 2), and letter (Level 1) to attend to. The selection is guided by considering the information gain of lexical elements at all levels, see Section 1 for further details.
@@
Figure 1(b) visually represents the saccades generated by our model while reading an example sentence: Active inference model of eye movements. The model demonstrates uncertainty about the first word (active) and therefore reads it using two saccades to the first (a) and fourth (i) letters. It follows the same process for the subsequent two words (inference and model). However, at this stage, it skips the next two words, as it perceives the most informative word to be the one following (movements). This mechanism showcases the model’s ability to dynamically select the most relevant information during the reading process.
@@
Figure 1(c) illustrates the probabilities of correctly recognized syllables (blue), words (red), and sentences (yellow) during the reading process. Over time, these probabilities approach one, indicating successful recognition. Additionally, the figure demonstrates that the levels operate at different time scales, with lower levels processing information more quickly than higher levels. This temporal separation arises because each level sends messages to the level above only after accumulating sufficient confidence (as shown by the vertical bars in the figure), a process that typically requires multiple rounds of inference (Friston et al., 2017).

### v3

@@ Abstract
We present a novel computational model employing hierarchical active inference to simulate reading and eye movements. The model characterizes linguistic processing as inference over a hierarchical generative model, facilitating predictions and inferences at various levels of granularity, from syllables to sentences. Our approach combines the strengths of large language models for realistic textual predictions and active inference for guiding eye movements to informative textual information, enabling the testing of predictions. The model exhibits proficiency in reading both known and unknown words and sentences, adhering to the distinction between lexical and nonlexical routes in dual route theories of reading. Our model therefore provides a novel approach to understand the cognitive processes underlying reading and eye movements, within a predictive processing framework. Furthermore, our model can potentially aid in understanding how maladaptive predictive processing can produce reading deficits associated with dyslexia. As a proof of concept, we show that attenuating the contribution of priors during the reading process leads to incorrect inferences and a more fragmented reading style, characterized by a greater number of shorter saccades, aligning with empirical findings regarding eye movements in dyslexic individuals. In summary, our model represents a significant advancement in comprehending the cognitive processes involved in reading and eye movements, with potential implications for understanding dyslexia in terms of maladaptive inference.
@@
Keywords: Hierarchical active inference, Predictive coding, Large Language Models, Reading, Dyslexia
1 Introduction
@@
Processing natural language – encompassing understanding, reading, and producing linguistic content – represents a fundamental ability of our species. Extensive research in psychology, neuroscience, linguistics, and machine learning has explored the intricate ways we process natural language. Neuroscientific studies have revealed that natural language processing is inherently hierarchical, involving multiple brain regions and the integration of various sensory inputs (capek2004cortical, ; butterfuss2018role, ; price2012review, ). This hierarchical processing spans from individual letters, phonemes, and words to complete sentence comprehension, with a hierarchy of brain areas actively maintaining these elements in working memory across multiple time scales (hasson2015hierarchical, ).
@@
Recent studies have increasingly supported the idea of hierarchical predictive coding as a formal theory describing perception as an inferential process, involving reciprocal exchanges between predictions and prediction errors across brain hierarchies (rao1999predictive, ; friston2005theory, ). In particular, various computational neuroimaging studies employing large language models (LLMs) during linguistic tasks have provided compelling evidence for both the predictive nature of language processing and the prediction hierarchies proposed by predictive coding (goldstein2022shared, ; schmitt2021predicting, ; baldassano2017discovering, ; zadbood2017we, ; chang2022information, ; caucheteux2023evidence, ; heilbron2022hierarchy, ). A recent computational neuroimaging study investigated human electrocorticographic (ECoG) responses to narratives using LLMs trained to predict the next word. Remarkably, the study revealed that like the LLMs, the brain engages in next-word prediction before word onset, computes prediction error signals, and utilizes latent representations of words (embeddings) contextualized based on the sequence of prior words (goldstein2022shared, ).
@@
Converging evidence emerges from three recent fMRI studies utilizing deep learning models during linguistic tasks. These studies not only confirm the predictive nature of language processing but also lend support to the prediction hierarchies proposed by predictive coding. The first fMRI study (schmitt2021predicting, ) trained a LLM to predict the next word at multiple time scales, identifying event boundaries as high surprise (also explored in (baldassano2017discovering, ; zadbood2017we, ; chang2022information, )). Analyzing human functional magnetic resonance data during story listening, the study revealed an event-based hierarchy of surprise signals evolving along temporoparietal regions, with surprise signals gating bottom-up and top-down connectivity across neighboring time scales. Furthermore, another fMRI study provided compelling evidence of the brain’s ability to predict a hierarchy of representations spanning multiple timescales in the future (caucheteux2023evidence, ). Enhancing LLMs with the capability to predict beyond the next word increased their fit with human data. The study also highlighted the hierarchical organization of brain predictions, ranging from temporal cortices predicting shorter-range representations (e.g., the next word) to frontoparietal cortices predicting higher-level, longer-range, and more contextual representations.
@@
Notably, a separate fMRI study reported that evoked brain responses to words are influenced by linguistic predictions and a metric of unexpectedness, closely aligning with the hierarchical predictive processing schemes, where lower-level predictions are informed by higher-level predictions (heilbron2022hierarchy, ). Additionally, the study demonstrated that these hierarchical predictions can be well-aligned with standard levels of analysis in psycholinguistics, including meaning, grammar, words, and speech sounds, reinforcing the validity of the standard decomposition. Taken together, these studies, alongside others (kutas1984brain, ; weissbart2020cortical, ; koskinen2020brain, ; willems2016prediction, ; frank2015erp, ), provide compelling support for the significance of prediction and hierarchical predictive coding in language processing. Moreover, they underscore the growing relevance of LLMs in comprehending linguistic processing (goldstein2022shared, ).
@@
Despite these advancements, the above studies have primarily focused on LLMs that passively receive sensory information, rather than actively searching for it. However, linguistic tasks, such as reading written text and listening to speech, are inherently active processes (ferro2010reading, ; norris2006bayesian, ; friston2021active, ). For example, during reading, eye movements (saccades) actively guide attention to relevant parts of the text, rather than processing every piece of text linearly. This active reading process suggests that saccades play a crucial role in hypothesis testing, selecting informative parts of the text to test predictions (ferro2010reading, ; norris2006bayesian, ; donnarumma2017action, ; friston2012perceptions, ).
@@
Several reading models incorporating eye movements, such as E-Z Reader (reichle1998toward, ), SWIFT (engbert2005swift, ; seelig2020bayesian, ; rabe2021bayesian, ), Über-Reader (reichle2021computational, ; veldre2020towards, ), Glenmore (reilly2002glenmore, ; reilly2006some, ), SEAM (rabe2024seam, ), the rational model of eye movements (bicknell2010rational, ) and OB1-Reader (snell2018ob1, ) significantly advanced our understanding of the role of reading dynamics. These models have successfully replicated numerous empirical findings, highlighting how word-level attributes like length, frequency, and predictability impact reading dynamics. However, these models do not exploit the generative capabilities of recent large language models and do not fully align with the aforementioned evidence of hierarchical predictive processing supporting reading and eye movements.
@@
In this paper, we propose a novel computational model that unifies hierarchical predictive processing and hypothesis testing during reading by integrating the LLM BERT (Bidirectional Encoder Representations from Transformers) (Devlin20194171, ) with active inference (friston2010free, ; parr2022active, ): a theory of perception and action based on Bayesian principles, whereby agents minimize expected surprise (or variational free energy) through iterative belief updating and action selection. Our model views reading as an active (Bayesian) inference problem, employing a hierarchical generative model to represent causal relationships between textual elements at different levels (letters, syllables, words, and sentences). By generating predictions at each level and testing them through saccades, our model actively simulates reading.
@@
The model incorporates three significant insights. Firstly, it conceptualizes linguistic processing as inference, employing a hierarchical generative model that allows for predicting and inferring at different timescales, such as syllables, words, and sentences. By integrating the LLM BERT (Devlin20194171, ) at the highest hierarchical level, our model can handle realistic reading tasks effectively, processing sentences of arbitrary length. Additionally, thanks to its hierarchical structure, our model can read both known words word-by-word and unknown words syllable-by-syllable, akin to lexical and nonlexical routes in dual route theories of reading (coltheart2005modeling, ).
@@
Secondly, the model utilizes its generative capabilities not only for language recognition and prediction, similar to LLMs, but also for simulating eye movements and saccades. In our model, reading involves an active, hypothesis testing process: the model generates saccades to the most informative parts of the text to validate its predictions and disambiguate among competing hypotheses about the content being read (ferro2010reading, ; baldassano2017discovering, ; donnarumma2017action, ; donnarumma2017you, ; friston2012perceptions, ; norris2006bayesian, ).
Thirdly, our model can be potentially used to gain insights on reading deficits, such as dyslexia. Dyslexia, a common reading disorder affecting 5-10% of the population, is associated not only with atypical brain activation patterns during language processing (price2012review, ) but also with atypical eye movement patterns, such as increased numbers of forward saccades and decreased saccade lengths compared to control groups (zoccolotti2005word, ; de1999eye, ; lyon2003definition, ; franzen2021individuals, ; hutzler2004eye, ; rayner1983eye, ). Crucially, it has been suggested that dyslexia could be characterized as a disorder of inference: a computational difficulty in effectively combining prior information, such as implicit memory of previous words, with noisy observations like the currently perceived word (jaffe2015computational, ). Specifically, the proposal is that the reading difficulties might arise because prior information is assigned excessively low weighting (or precision) relative to its internal noise. The low precision-weighting increases the time required to form a coherent understanding of text (or to perform auditory discrimination, as in (jaffe2015computational, )), because novel observations consistently surprise dyslexics. By incorporating this proposal in our model, we provide a proof of concept that it can qualitatively reproduce eye movement patterns associated with dyslexia.
@@ 2 Methods
In this section, we provide a summary description of the model, a brief introduction to active inference (parr2022active, ), and a detailed discussion of the hierarchical active inference reading model used in this paper.
@@ 2.1 Brief explanation of the model
We present a novel hierarchical active inference model for reading and eye movements during reading (friston2010free, ). Consistent with evidence of hierarchical language processing in the brain (heilbron2022hierarchy, ), our model consists of three levels representing syllables, words, and sentences, as illustrated in Figure 1(a-b). At each level, the inference process involves integrating three types of messages: bottom-up messages from nodes at the level below (conveying observations), top-down messages from nodes at the level above (conveying predictions), and lateral messages from nodes representing the previous timestep at the same level (providing memory). For instance, the inference about the current syllable is informed by bottom-up observations about the currently observed letter, top-down predictions from the currently inferred word, and lateral information about the previous syllable. Similarly, the inference about the current word relies on bottom-up observations (i.e., the currently estimated syllable), top-down predictions from the level above (i.e., at the sentence level), and lateral information about the previous word.
@@
Figure 1: The hierarchical active inference model of reading and eye movements. (a) The hierarchical model comprises three layers for syllables (Level 1), words (Level 2), and sentences (Level 3), represented using the formalism of Partially Observable Markov Decision Processes (POMDPs). Empty nodes indicate hidden variables, filled nodes represent observations, and edges depict probabilistic relations between the nodes, as detailed in B. (b) Model parameters. (c) The sequence of saccades generated by the model while reading an example sentence. Yellow dots indicate the positions of the saccades in the text, and colored lines trace the sequence of saccades from red to yellow. (d) Evolution of the probabilities of correctly recognized syllables (blue lines), words (orange lines), and sentences (yellow lines) over time, while reading the sentence of Panel c. The blue circles at the bottom level indicate saccades. Once the first level confidently infers a syllable (or when the maximum level of iterations is reached), it sends a bottom-up signal to the second level (indicated by the orange vertical bar) to aid in word inference. Similarly, the second level sends a signal to the third level (indicated by the yellow vertical bar) upon confidently inferring a word. Please refer to the main text for further explanation.
To ensure that our model is capable of reading realistic text, we incorporate the large language model BERT (Devlin20194171, ) to generate prior probability distributions over the current word when reading single words, and over the next words based on previously inferred words when reading sentences. BERT was chosen because it offers computationally efficient access to explicit probability distributions over upcoming words and sentence continuations. Given an initial context—a sequence of already read words—and a predictive horizon of $H$ steps, BERT produces $K$ candidate sentence completions. These candidates are used to construct nested multinomial distributions over sentences, their constituent words, and the syllables comprising those words. These distributions then serve as priors for the three levels of the hierarchical active inference model, thereby initializing the inference process to estimate the most plausible continuation of the sentence. Inference proceeds over successive time steps as new words are observed. After a fixed number of inference steps, BERT is queried again with the updated context—including the newly read words—and the process is repeated. This iterative procedure enables the model to read and infer sentences of varying lengths.
@@
Lastly, our model has the capability to determine where to make (forward or backward) saccades and select the next letter observation (donnarumma2017action, ). The process of selecting where to direct the next saccade (i.e., which position in the syllable) is guided by the expected information gain of lexical elements across all three levels of the model. Specifically, saccades are directed to the syllable position where the resulting letter observation is expected to reduce uncertainty about the current syllable (Level 1), word (Level 2), and sentence (Level 3); see Section 2 for further details.
@@
Figure 1(c) visually represents the saccades generated by our model while reading an example sentence: Active inference model of eye movements. The model demonstrates uncertainty about the first word (active) and therefore reads it using two saccades to the first (a) and fourth (i) letters. It follows the same process for the subsequent two words (inference and model). However, at this stage, it skips the next two words, as it perceives the most informative word to be the one following (movements). This mechanism showcases the model’s ability to dynamically select the most relevant information during the reading process.
@@
Figure 1(d) illustrates the probabilities of correctly recognized syllables (blue), words (red), and sentences (yellow) during the reading process. Over time, these probabilities approach one, indicating successful recognition. Additionally, the figure demonstrates that the levels operate at different time scales, with lower levels processing information more quickly than higher levels. This temporal separation arises because each level sends messages to the level above only after accumulating sufficient confidence (as shown by the vertical bars in the figure), a process that typically requires multiple rounds of inference (friston2017active, ). Note that the time it takes to arrive at a high probability of syllable can be interpreted as saccade duration – implying that the longer the time required to resolve uncertainty about the fixated syllable, the longer the saccade duration.